OpenStudy (anonymous):

The formula a=d/g gives the average gas mileage a of a vehicle that uses g gallons of gas to travel d miles. Use the formula to find how many gallons of gas a vehicle with an average gas mileage of 20.2 miles per gallon will use to travel 75 miles. Round to nearest tenth.

OpenStudy (anonymous):

1)s=n/2(a+l) distribute the 2 s=n/2a+2l mutiply by (2a+2l) s(2a+2l)=n distribute the s 2as+2al=n subtract 2as 2al=n-2as divide by 2a l=(n-2as)/2a 2) a=d/g 20.2=75/d multiply by d 20.2d=75 divide by 20.2 d=3.7 miles

OpenStudy (anonymous):

20.2=75/g g=75/20.2 g=3.71 gallons